a batsman has average of 20 runs in the first six tests. in the second to the seventh tests his average score was 21. if he scored a total of 30 runs in the first and sevenths test.find the score in seventh test.

Answer by Boreal(15235)

x=first test
30-x=seventh test
Total for the first 6 tests is average *n=120
Total for the second-seventh tests is 126
second through sixth tests=120-x
second through seventh tests=126; the seventh test was therefore 6+x
The seventh test, however is also 30-x
Therefore 30-x=6+x; 2x=24, x=12
I get 30-x=18.
Check: First is 12, so second through six would be 108. Seventh is 18 for a total of 126.
Test 16: First is 14, so second through six would be 106. Seventh is 16 for a total of 122.
If the answer is 16, I apologize, but I don't see it. Somebody else will.